# Stabilizing Rate of Stochastic Control Systems with Multiplicative Noise

### v1

@@ 5.2 Certified Performance Bounds
For each $\tau\in(0,1)$ the fixed point $P^{(\tau)}$ satisfies
$\gamma^{(\tau)}\ \leq\ (1-\tau)C_{A}+\tau.$
Therefore the upper bound in (25) can be further relaxed to
$J^{*}\ \leq\ \log\!\frac{(1-\tau)C_{A}+\tau}{1-\tau}.$
Unlike the original expression, this relaxed bound is explicit and strictly decreases as $\tau\downarrow 0$. Thus, even without monotonicity of the original upper estimates, one obtains a monotone and computable envelope bounding them from above. The parameter $\tau$ acts as a regularizer: it guarantees the existence of interior fixed points $P^{(\tau)}\succ 0$ and produces computable two-sided bounds in (25).
@@ 5.3 Regularized Normalized Value Iteration
    RNVI converges globally and linearly to $P^{(\tau)}$:
@@
The condition $\Lambda(\tau)<1$ is a sufficient (but not necessary) criterion. In particular, since $\Lambda(\tau)\to 0$ as $\tau\uparrow 1$, there always exists $\tau_{0}$ close to $1$ such that $\Lambda(\tau_{0})<1$. If $\Lambda(\tau)\geq 1$ and global contraction cannot be certified, one can still reach the desired fixed point by a continuation procedure with respect to $\tau$, as detailed in Theorem 13.
@@
an affine function of $H$ whose coefficients depend continuously on $P$. Thus $(P,H)\mapsto D\Phi(P)[H]$ is continuous.
@@
1:Input: System matrices $(A,\bar{A},B,\bar{B})$; noise $\omega_{k}\sim\mathcal{N}(0,\sigma^{2})$; initial $P_{0}\succ 0$ with $\operatorname{Tr}(P_{0})=1$; decreasing schedule $\tau_{0}>\tau_{1}>\cdots>\tau_{M}=\tau_{\mathrm{target}}$; tolerance $\varepsilon$.
@@
Suppose $\tau_{0}\in(0,1)$ satisfies $\Lambda(\tau_{0})<1$, and let $\tau\in(0,\tau_{0})$ be a target parameter. Then, there exists an integer $M\geq 1$ and a decreasing sequence $\tau_{0}>\tau_{1}>\cdots>\tau_{M}=\tau$ such that, for each $j=0,\dots,M-1$, RNVI with parameter $\tau_{j+1}$, initialized at $P^{(\tau_{j})}$, converges locally to $P^{(\tau_{j+1})}$. In particular, this continuation procedure yields convergence to the fixed point $P^{(\tau)}$ at the target $\tau$.
@@
Let $\tau_{0}\in(0,1)$ be such that $\Lambda(\tau_{0})<1$, and fix a target $\tau\in(0,\tau_{0})$. By Lemma 5, the local Lipschitz modulus ${\rm Lip}(P,\tau)$ depends continuously on $(P,\tau)$ on $S_{a}\times(0,1)$. We construct the schedule $\{\tau_{j}\}$ inductively. For the case $j=0$, note that ${\rm Lip}(P^{(\tau_{0})},\tau_{0})\ <\ 1.$ At step $j$, assume that $\tau_{j}\in(0,1)$ has been chosen and that ${\rm Lip}(P^{(\tau_{j})},\tau_{j})<1.$ By the continuity of ${\rm Lip}(P,\tau)$ at $(P^{(\tau_{j})},\tau_{j})$, there exists $\Delta\tau_{j}>0$ such that for all $\tilde{\tau}\in(\tau_{j}-\Delta\tau_{j},\tau_{j}]$ , ${\rm Lip}(P^{(\tau_{j})},\tilde{\tau})<1$ holds and hence $\widehat{\Phi}_{\tilde{\tau}}$ is a contraction in a neighborhood of $P^{(\tau_{j})}$ for all such $\tilde{\tau}$.
@@
$\tau_{j+1}\in\bigl(\max\{\tau,\tau_{j}-\Delta\tau_{j}\},\ \tau_{j}\bigr),$
which is always possible as long as $\tau_{j}>\tau$. Since $\delta_{\tau}$ is strictly increasing in $\tau$, we have $\mathcal{S}_{\delta_{\tau_{j}}}\subset\mathcal{S}_{\delta_{\tau_{j+1}}}$ whenever $\tau_{j+1}<\tau_{j}$; hence $P^{(\tau_{j})}\in\mathcal{S}_{\delta_{\tau_{j+1}}}$, it is a valid initialization point for the RNVI iteration at parameter $\tau_{j+1}$.
Because $\widehat{\Phi}_{\tau_{j+1}}$ is a contraction around $P^{(\tau_{j})}$, RNVI with parameter $\tau_{j+1}$, initialized at $P^{(\tau_{j})}$, converges locally to its fixed point $P^{(\tau_{j+1})}$. This implies
@@
and the induction continues. Thus one can construct a decreasing sequence $\{\tau_{j}\}$ approaching the target $\tau$, and RNVI converges locally at each stage. ∎
@@
This continuation procedure is practically robust: start at a “safe” $\tau_{0}$ where $\Lambda(\tau_{0})<1$, then decrease $\tau$ in small steps, reusing the previous fixed point as initialization. Each step enjoys local contraction, and the process converges to $P^{(\tau)}$ for the desired target $\tau$.
@@
Relation to Riccati theory and numerical methods. The equation (23) may be viewed, at an algebraic level, as a generalized Riccati-type nonlinear matrix equation on $\mathbb{S}_{++}^{n}$. The operator $\Phi$ has a structural form similar to the generalized Riccati operator that appears in discrete-time stochastic LQ problems with multiplicative noise. In this broad sense, (23) is Riccati-type, though its analytical role differs fundamentally from that of classic LQ Riccati equations.
Classical Riccati theory and numerical methods. In continuous and discrete-time LQ control, algebraic Riccati equations (AREs) characterize optimal regulators and stabilizing feedback gains under stabilizability and detectability assumptions; see the standard references (Anderson and Moore (2007); Bittanti et al. (1991)). These structural properties underpin several major families of numerical methods: (i) Hamiltonian/symplectic approaches that compute a stable invariant subspace of the associated Hamiltonian matrix Laub (1979); (ii) Newton-Kleinman iterations, which solve a Lyapunov equation at each step (Feitzinger et al. (2009)); and (iii) projection and Krylov-based large-scale solvers (Bunse-Gerstner (1996)). The convergence of these algorithms is typically explained through Loewner-order monotonicity, extremal characterizations of the stabilizing solution, and Hamiltonian invariant subspace geometry.
Why do these methods not apply directly in our setting? Although $\Phi$ has a Riccati-type algebraic pattern, the present problem is not an LQ control problem but a Bellman-type nonlinear eigenvalue problem $\Phi(P)=\gamma P$, whose solution determines the optimal stabilizing rate. This eigen-problem is homogeneous and scale invariant, and the regularization in (23) is essential. Consequently: (i) $\widehat{\Phi}_{\tau}(P)=P$ is not a standard ARE and does not correspond to a stabilizing closed-loop linear system; (ii) there is no Hamiltonian matrix, so Hamiltonian/symplectic solvers do not apply; and (iii) trace-normalization destroys Loewner-order monotonicity, preventing the use of monotone Riccati iterations. For these reasons, the contraction framework in Theorem 12 is the natural tool for establishing the uniqueness and algorithmic convergence.
Future directions. The order structure of $P\mapsto\Phi(P)$ suggests possible connections to nonlinear Perron–Frobenius theory and Hilbert’s projective metric (Lemmens and Nussbaum (2012)). Developing cone-theoretic uniqueness results or dynamical systems interpretations (e.g., generalized Riccati flows) may lead to stronger analytical guarantees beyond contractivity.
@@ 6.1 Two-dimensional system: noise-level sweep
Methodology.
Following the construction in Section 5, each RNVI fixed point $P^{(\tau)}$ provides computable bounds on the asymptotic growth rate:
@@
Since $J^{*}=2\log\rho^{*}$, these bounds immediately yield an interval for $\rho^{*}$.
@@
Results.
Figure 1 shows, for each noise level $\sigma$, the lower (blue) and upper (red) bounds for $\rho^{*}$ across the $\tau$-grid which decreasing to the right. Table 1 then summarizes the best bounds across $\tau$, and reports the corresponding intervals for both $J^{*}$ and $\rho^{*}$.
@@ Results.
Discussion.
The results highlight two consistent features. First, as the noise intensity $\sigma$ increases, the certified optimal stabilizing rate $\rho^{*}$ moves closer to 1, reflecting the expected degradation of mean-square stability under stronger multiplicative disturbances. Second, the certified gaps remain very small, on the order of $10^{-3}\sim 10^{-2}$ for $\rho^{*}$. This tightness demonstrates that RNVI yields not only reliable existence and convergence guarantees but also numerically precise performance bounds across a wide range of noise levels. For illustration, we also inspect the closed-loop behavior at $\sigma=2$ using the RNVI-derived approximate optimal gain. A Monte Carlo simulation shows that $\mathbb{E}[x_{k}^{\top}x_{k}]$ decays nearly linearly on a logarithmic scale, indicating clear exponential mean-square stabilization under this feedback.

### v2

@@ 5.2 Certified Performance Bounds
Having obtained computable upper and lower bounds for $\rho^{*}$ in Theorem 11, we next analyze how tight these bounds are as $\tau$ varies.
Theorem 12.
For any $\tau\in(0,1)$ and any solution $P^{(\tau)}$ of (23), define
$L_{\tau}:=L\big(P^{(\tau)}\big),\quad U_{\tau}:=\frac{\gamma^{(\tau)}}{1-\tau},\quad\Delta_{\tau}:=U_{\tau}-L_{\tau}.$
Then, the following statements hold.
(i) The gap between the squared certified bounds for $\rho^{*}$ in Theorem 11 admits the explicit expression
$\Delta_{\tau}=\frac{\tau}{n(1-\tau)}\,\lambda_{\max}\big((P^{(\tau)})^{-1}\big).$
(ii) Let $\{\tau_{k}\}_{k\geq 1}\subset(0,1)$ with $\tau_{k}\downarrow 0$, and let $P^{*}\succeq 0$ be a limit point of $\{P^{(\tau_{k})}\}$ as characterized in Theorem 8. If the limit satisfies $P^{*}\succ 0$, then the certified squared bound becomes asymptotically tight along the corresponding convergent subsequence:
$\Delta_{\tau_{k}}\ \longrightarrow\ 0\qquad\text{as }k\to\infty.$
Proof.
Fix $\tau\in(0,1)$ and a corresponding fixed point $P^{(\tau)}\succ 0$ of (23). By the fixed-point relation (23), it holds
$(1-\tau)\,\Phi(P^{(\tau)})+\tfrac{\tau}{n}I=\gamma^{(\tau)}P^{(\tau)}.$
Multiplying by $({P^{(\tau)}})^{-1/2}$ on both sides of above equation and rearranging it yields
$(1-\tau)\,({P^{(\tau)}})^{-1/2}\Phi(P^{(\tau)})({P^{(\tau)}})^{-1/2}=\gamma^{(\tau)}I-\tfrac{\tau}{n}(P^{(\tau)})^{-1}.$
Hence, the eigenvalues of $({P^{(\tau)}})^{-1/2}\Phi(P^{(\tau)})({P^{(\tau)}})^{-1/2}$ are
$\lambda_{i}=\frac{\gamma^{(\tau)}}{1-\tau}-\frac{\tau}{n(1-\tau)}\,\mu_{i},\quad i=1,\cdots,n,$
where $\mu_{i}$ is the corresponding eigenvalues of $(P^{(\tau)})^{-1}$. Therefore, it holds
$L_{\tau}=\min_{i}\lambda_{i}=\frac{\gamma^{(\tau)}}{1-\tau}-\frac{\tau}{n(1-\tau)}\,\lambda_{\max}((P^{(\tau)})^{-1}),$
which proves part (i).
For part (ii), if $P^{*}\succ 0$, then $\lambda_{\min}(P^{(\tau_{k})})\to\lambda_{\min}(P^{*})>0$ and hence $\lambda_{\max}\big((P^{(\tau_{k})})^{-1}\big)$ stays bounded. Since $\tau_{k}/(1-\tau_{k})\to 0$, it holds
$\Delta_{\tau_{k}}=\frac{\tau_{k}}{n(1-\tau_{k})}\,\lambda_{\max}\big((P^{(\tau_{k})})^{-1}\big)\longrightarrow 0.$
We complete the proof. ∎
@@
The condition that the subsequential limit $P^{*}$ satisfies $P^{*}\succ 0$ in Theorem 12 is mathematically strong, but it can be relaxed. Indeed, the tightness requirement there only involves the product
$\frac{\tau}{\,1-\tau\,}\,\lambda_{\max}\bigl((P^{(\tau)})^{-1}\bigr),$
and asymptotic tightness follows whenever this quantity vanishes as $\tau\downarrow 0$, namely,
$\lim_{\tau\downarrow 0}\frac{\tau}{\,1-\tau\,}\,\lambda_{\max}\bigl((P^{(\tau)})^{-1}\bigr)=0.$
Thus, strict positivity of the limiting matrix $P^{*}$ is not required. It suffices that $\lambda_{\max}\bigl((P^{(\tau)})^{-1}\bigr)$ grows more slowly than $1/\tau$ as $\tau\downarrow 0$, such that the decay of $\tau$ dominates the growth of $\lambda_{\max}\bigl((P^{(\tau)})^{-1}\bigr)$. In general, this relaxed condition is mild: for sufficiently small $\tau$, the product $\tau\,\lambda_{\max}\bigl((P^{(\tau)})^{-1}\bigr)$ can be numerically small in practice. Actually, as documented in Section 6, this relaxed condition is satisfied across all 25 diverse numerical experiments reported in this paper.
@@ 5.3 Regularized Normalized Value Iteration
    The sequence $(P_{k})_{k\in\mathbb{N}}$ generated by RNVI converges globally and linearly to $P^{(\tau)}$:
@@
The condition $\Lambda(\tau)<1$ is a sufficient (but not necessary) criterion. In particular, since $\Lambda(\tau)\to 0$ as $\tau\uparrow 1$, there always exists $\tau_{0}$ close to $1$ such that $\Lambda(\tau_{0})<1$.
@@
an affine function of $H$ whose coefficients depend continuously on $P$. Thus, $(P,H)\mapsto D\Phi(P)[H]$ is continuous.
@@
1:Input: System matrices $(A,\bar{A},B,\bar{B})$; noise $\omega_{k}\sim\mathcal{N}(0,\sigma^{2})$; initial $P_{0}\succ 0$ with $\operatorname{Tr}(P_{0})=1$; decreasing schedule $\tau_{0}>\tau_{1}>\cdots>\tau_{M}$; tolerance $\varepsilon$.
@@
Suppose $\tau_{0}\in(0,1)$ satisfies $\Lambda(\tau_{0})<1$. Then, one can construct a strictly decreasing sequence of parameters
$\tau_{0}>\tau_{1}>\tau_{2}>\cdots,$
together with associated fixed points $\{P^{(\tau_{j})}\}_{j\geq 0}$ of (23), such that at each stage the sequence generated by RNVI with parameter $\tau_{j+1}$, initialized at $P^{(\tau_{j})}$, converges locally to $P^{(\tau_{j+1})}$. In particular, the continuation procedure produces a decreasing family of parameters $\{\tau_{j}\}$ and fixed points $\{P^{(\tau_{j})}\}$ linked by local RNVI convergence.
@@
Let $\tau_{0}\in(0,1)$ be such that $\Lambda(\tau_{0})<1$, and fix a target $\tau\in(0,\tau_{0})$. By Lemma 5, the local Lipschitz modulus ${\rm Lip}(P,\tau)$ depends continuously on $(P,\tau)$ on $S_{a}\times(0,1)$. We construct a decreasing sequence $\{\tau_{j}\}$ inductively. For $j=0$, note that
${\rm Lip}(P^{(\tau_{0})},\tau_{0})<1.$
At step $j$, suppose $\tau_{j}$ has been chosen and
${\rm Lip}(P^{(\tau_{j})},\tau_{j})<1$
holds. By the continuity of ${\rm Lip}(P,\tau)$ at $(P^{(\tau_{j})},\tau_{j})$, there exists $\Delta\tau_{j}>0$ such that for every $\tilde{\tau}\in(\tau_{j}-\Delta\tau_{j},\tau_{j}]$,
${\rm Lip}(P^{(\tau_{j})},\tilde{\tau})<1$
holds and consequently $\widehat{\Phi}_{\tilde{\tau}}$ is a contraction in a neighborhood of $P^{(\tau_{j})}$.
@@
$\tau_{j+1}\in\bigl(\max\{\tau,\tau_{j}-\Delta\tau_{j}\},\,\tau_{j}\bigr).$
Since $\delta_{\tau}$ is strictly increasing in $\tau$, we have $\mathcal{S}_{\delta_{\tau_{j}}}\subset\mathcal{S}_{\delta_{\tau_{j+1}}}$ whenever $\tau_{j+1}<\tau_{j}$. Hence $P^{(\tau_{j})}\in\mathcal{S}_{\delta_{\tau_{j+1}}}$ and is a valid initialization for RNVI at parameter $\tau_{j+1}$.
Because $\widehat{\Phi}_{\tau_{j+1}}$ is a contraction near $P^{(\tau_{j})}$, the sequence generated by RNVI with parameter $\tau_{j+1}$, initialized at $P^{(\tau_{j})}$, converges locally to the fixed point $P^{(\tau_{j+1})}$. This yields
@@
allowing the induction to proceed.
In this manner, one obtains a decreasing sequence $\{\tau_{j}\}$ with local convergence at each stage, and the continuation procedure produces the fixed point $P^{(\tau)}$. ∎
@@
The fixed point of (23) is known to exist by Brouwer’s fixed-point theorem, but that theorem is nonconstructive: it does not explain how to obtain an actual fixed point in practice. The continuation result in Theorem 14 addresses this gap by providing a problem-specific mechanism that produces a decreasing family $\{\tau_{j}\}$ together with associated fixed points $\{P^{(\tau_{j})}\}$, each obtained by locally convergent RNVI iteration starting from the preceding stage. Its limitation is that it does not guarantee that an arbitrarily prescribed target value of $\tau$ can be reached along such a path.
As an alternative, one may resort to constructive methods based on Sperner-type simplicial subdivision for a continuous self-map $F:K\to K$ on a compact convex set. By refining a simplicial mesh of $K$ and labeling each vertex according to the direction of F(x)xF(x)-x, Sperner’s lemma ensures the existence of a fully labelled small simplex; points within this simplex provide approximate fixed points with accuracy proportional to the mesh size.
@@
Relation to Riccati theory and numerical methods. Equation (23) can be viewed algebraically as a Riccati-type nonlinear matrix equation on $\mathbb{S}_{++}^{n}$, since $\Phi$ has a structure analogous to generalized Riccati operators in discrete-time stochastic LQ problems with multiplicative noise. Nevertheless, its role here differs from classical LQ Riccati equations.
Classical Riccati theory and numerical methods. In continuous- and discrete-time LQ control, algebraic Riccati equations (AREs) characterize optimal regulators and stabilizing feedback gains under standard stabilizability/detectability assumptions; see Anderson and Moore (2007); Bittanti et al. (1991). This structure supports, among others: (i) Hamiltonian/symplectic methods based on stable invariant subspaces Laub (1979); (ii) Newton–Kleinman iterations solving a Lyapunov equation at each step Feitzinger et al. (2009); and (iii) projection/Krylov large-scale solvers Bunse-Gerstner (1996).
Why do these methods not apply directly in our setting? Although $\Phi$ is Riccati-type in form, our problem is a Bellman-type nonlinear eigenvalue problem $\Phi(P)=\gamma P$, and the regularization in (23) is essential. Accordingly, $\widehat{\Phi}_{\tau}(P)=P$ is not a standard ARE; there is no associated Hamiltonian matrix for symplectic solvers; and trace normalization breaks the Loewner-order monotonicity used in many Riccati iterations. Hence we instead rely on the contraction framework in Theorem 13, which provides a direct and rigorous basis for RNVI.
Future directions. The order structure of $P\mapsto\Phi(P)$ suggests possible links to nonlinear Perron–Frobenius theory and Hilbert’s projective metric Lemmens and Nussbaum (2012). Developing cone-theoretic uniqueness results or dynamical-systems interpretations (e.g., generalized Riccati flows) may yield stronger guarantees beyond contractivity.
@@ 6.1 Two-dimensional system: noise-level sweep
Methodology. Following the construction in Section 5, each RNVI fixed point $P^{(\tau)}$ provides computable bounds on the asymptotic growth rate:
@@
Since $J^{*}=2\log\rho^{*}$, these bounds immediately yield an interval for $\rho^{*}$. Theorem 12 shows that the tightness of the RNVI certificates is governed by the squared gap
$\Delta_{\tau}=\frac{\tau}{n(1-\tau)}\,\lambda_{\max}\!\bigl(P^{(\tau)\,-1}\bigr).$
This motivates a numerical inspection of how $\Delta_{\tau}$ behaves as $\tau\downarrow 0$, and whether the matrices $P^{(\tau)}$ remain well conditioned.
Motivated by Theorem 12 and Remark Remark, we therefore include two additional diagnostic plots: (i) the squared-gap curve $\Delta_{\tau}$, and (ii) the conditioning measure $\lambda_{\max}\bigl((P^{(\tau)})^{-1}\bigr)=1/\lambda_{\min}(P^{(\tau)})$, both plotted across $\tau$ for all noise levels $\sigma$. The first plot directly illustrates the decay of the certified squared gap as $\tau$ decreases, while the second provides a numerical indication of whether the relaxed tightness condition in Remark Remark is plausibly satisfied over the explored range of $\tau$.
@@
In addition, for each noise level $\sigma$ we also record the number of RNVI iterations required for convergence at every value of $\tau$.
Results. Figure 1 shows, for each noise level $\sigma$, the lower (blue) and upper (red) bounds for $\rho^{*}$ across the $\tau$-grid which decreasing to the right. Table 1 then summarizes the best bounds across $\tau$, and reports the corresponding intervals for both $J^{*}$ and $\rho^{*}$.
@@
Figure 2: Certified squared-gap $\Delta_{\tau}=\frac{\tau}{n(1-\tau)}\lambda_{\max}\big((P^{(\tau)})^{-1}\big)$ for each noise level $\sigma$.
Figure 3: $\lambda_{\max}\big((P^{(\tau)})^{-1}\big)=1/\lambda_{\min}(P^{(\tau)})$ for each noise level $\sigma$.
Figures 23 reveal two important and consistent phenomena:
  1. 1.
    The certified squared-gap $\Delta_{\tau}$ decreases rapidly as $\tau$ becomes small for all noise levels $\sigma$. Numerically, $\Delta_{\tau}$ is already below $10^{-2}$ once $\tau\approx 10^{-3}$, and its decay is consistent with the qualitative trend described by Theorem 12.
  2. 2.
    $\lambda_{\max}\big((P^{(\tau)})^{-1}\big)$ remains well bounded across $\sigma$ and across all small $\tau$. This behavior supports the relaxed condition stated in Remark Remark, indicating that no noticeable deterioration occurs as $\tau$ decreases.
Together, these two diagnostic plots confirm that RNVI produces well-conditioned fixed points near $\tau=0$ and that the certified bounds become extremely tight in this regime.
Figure 4: RNVI iteration counts for $\sigma=3$ across the $\tau$–grid.
Figure 4 reports the RNVI iteration count for $\sigma=3$ across the $\tau$–grid. Although the continuation Theorem 14 does not guarantee that an arbitrary preassigned sequence of $\tau$–values will always lie within the local basin of convergence, the observed behavior is highly stable: the iteration count remains uniformly small (between 12 and 30) over more than five orders of magnitude in $\tau$. Finally, we note that for every noise level and every value of $\tau$ in the sweep, the RNVI iteration converges in only a small number of steps. This rapid convergence is consistently observed not only in this noise–sweep experiment but also in all other numerical studies reported later in the paper. To avoid disrupting the flow, we do not tabulate all iteration counts, but the behavior illustrated in Figure 4 is representative of the overall performance.
@@
Discussion. The results highlight two consistent features. First, as the noise intensity $\sigma$ increases, the certified optimal stabilizing rate $\rho^{*}$ becomes larger, reflecting the expected degradation of mean-square stability under stronger multiplicative disturbances. Second, the certified gaps remain very small, on the order of $10^{-3}\sim 10^{-2}$ for $\rho^{*}$. This tightness demonstrates that RNVI yields not only reliable existence and convergence guarantees but also numerically precise performance bounds across a wide range of noise levels.
Figures 23 illustrate the behavior of the quantity $\tau\,\lambda_{\max}\big((P^{(\tau)})^{-1}\big)$ as $\tau\downarrow 0$. In all cases considered, this product remains small for sufficiently small $\tau$, leading to a rapid decay of the certified squared-gap $\Delta_{\tau}$. These observations are fully consistent with the relaxed tightness condition discussed in Remark Remark, and indicate that the sufficient condition in Theorem 12 (ii) is, at least numerically, not restrictive in the present setting.
For illustration, we also inspect the closed-loop behavior at $\sigma=2$ using the RNVI-derived approximate optimal gain. A Monte Carlo simulation shows that $\mathbb{E}[x_{k}^{\top}x_{k}]$ decays nearly linearly on a logarithmic scale, indicating clear exponential mean-square stabilization under this feedback.